OpenStudy (anonymous):

please find the product -x^3/x^2+5x-14 times ( 2-x)

OpenStudy (kropot72):

\[\frac{-x ^{3}}{x ^{2}+5x-14}\times (2-x)=\frac{x ^{3}(x-2)}{(x+7)(x-2)}=\frac{x ^{3}}{x+7}\]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)